# Distance-Adaptive Quaternion Knowledge Graph Embedding with Bidirectional Rotation

### v1

@@ Abstract
Quaternion contains one real part and three imaginary parts, which provided a more expressive hypercomplex space for learning knowledge graph. Existing quaternion embedding models measure the plausibility of a triplet either through semantic matching or geometric distance scoring functions. However, it appears that semantic matching diminishes the separability of entities, while the distance scoring function weakens the semantics of entities. To address this issue, we propose a novel quaternion knowledge graph embedding model. Our model combines semantic matching with entity’s geometric distance to better measure the plausibility of triplets. Specifically, in the quaternion space, we perform a right rotation on head entity and a reverse rotation on tail entity to learn rich semantic features. Then, we utilize distance adaptive translations to learn geometric distance between entities. Furthermore, we provide mathematical proofs to demonstrate our model can handle complex logical relationships. Extensive experimental results and analyses show our model significantly outperforms previous models on well-known knowledge graph completion benchmark datasets. Our code is available at https://github.com/llqy123/DaBR.
@@ 1 Introduction
Knowledge graphs (KGs) Liang et al. (2024a) are powerful tools for representing valid factual triplets by capturing entities and their relationships in a graphical format. Owing to the well-structured graph, KGs are often used for various Natural Language Processing tasks including but not limited to, question answering Mendes et al. (2024); Faldu et al. (2024), entity alignment Wang et al. (2024), KG-based recommendation Liang et al. (2024c) and KG enhanced Large Language Model Wen et al. (2024).
However, KGs are usually incomplete, and the incompleteness limits their application. As an effective tool for completing missing facts, knowledge graph completion (KGC) has received much attention from researchers. Typically, researchers transform KGC tasks into knowledge graph embeddings (KGEs). KGE refers to learning representations of entities and relations in a low-dimensional space while preserving graph’s inherent structure and semantic properties. In this representation space, a scoring function can be defined to measure the plausibility of each triplet, where valid triplets should receive higher scores than these invalid ones.
@@
Quaternion contains one real part and three imaginary parts, which providing a more expressive space for learning embeddings of entities and relations. Rotation in quaternion is often used to model the KGs. For example, QuatE Zhang et al. (2019) learns semantic information about entities by treating relations as rotations from head entities to tail entities. TransERR Li et al. (2024) encodes the KG by rotating the head and tail entities with their corresponding unit quaternions. These models use either semantic matching or distance scoring functions to measure the plausibility of the triplet, respectively. However, it appears that semantic matching diminishes the separability of entities, while the distance scoring function weakens the semantics of entities from our visualization analysis in Figure 1  [^1] For more information about queries, see Section 6.4.. Specifically, as shown in Figure 1, we can observe that QuatE using semantic matching as a scoring function overlaps in each queries. The entities of TransERR using the distance scoring function are also indistinguishable from each query.
To address this issue, we propose a Distance-adaptive quaternion knowledge graph embedding with Bidirectional Rotation model, named as DaBR. Our model combines semantic matching with entity’s geometric distance to better measure the plausibility of triplets. Specifically, in the quaternion space, we perform a right rotation on the head entity and a reverse rotation on the tail entity to learn the rich semantic features. This process is called bidirectional rotation. We conducted extensive experiments on multiple well-known benchmark datasets of knowledge graph completion task. The experimental results and analyses demonstrated the effectiveness and robustness of our model.
Our contribution is summarized as follows:
  • We propose to perform a right rotation on the head entity and a reverse rotation on the tail entity to learn the rich semantic features.
  • We propose to learn the embedding distance between entities by incorporating distance adaptive translations.
@@
  • Extensive experiments show that our model provides consistent and significant improvements over previous models on the vast majority of metrics.
@@ 2 Related Work
For KGE models, the design of the scoring function directly affects these models’ performance and effectiveness. Based on the calculation methods of scoring functions in previous models, KGE scoring functions can be mainly categorized into semantic matching- and geometric distance-based.
@@
Semantic matching. Semantic matching scoring functions capture the interactions between entities and relations through inner products on embedding vectors. The hypothesis is that entities connected by relations are close to each other in the semantic space. For example, QuatE Zhang et al. (2019) obtains entity semantic information through the Hamiltonian rotation of the head entity on the relation in quaternion space. DualE Cao et al. (2021) further enhances QuatE to model knowledge graphs in dual quaternion space. QuatRE Nguyen et al. (2022) associates each relation with two relation-aware rotations, which are used to rotate the quaternion embeddings of the head and tail entities, respectively.
@@
A common feature of these models is the computation of the inner product between the head entity and the tail entity after a relation transformation. However, these models overlook the geometric distance properties between entities in the knowledge graph, leading to distorted embeddings of the learned entities.
@@
Geometric distance. Geometric distance scoring functions assess the plausibility of triplets by calculating the distances between embedding vectors in the representation space. The goal of this scoring function is keep the head/tail entity vector closer to the tail/head entity vector after being transformed through the relation vector. For example, TransE Bordes et al. (2013), considered the first model to employ a geometric distance scoring function, assumes that triplets $(h,r,t)$ in knowledge graphs should satisfy the expression $h+r\approx t$. However, TransE struggles with more complex relation types, such as one to many (1-to-N), many to one (N-to-1) and many to many (N-to-N).
@@
To address this limitation, several models using distance-based scoring functions have been proposed. For example, Rotate3D Gao et al. (2020) maps entities to an 3D space, defining the relation as a rotation from the head entity to the tail entity. Trans4E Nayyeri et al. (2021) performs rotations and translations in a quaternion space. RotateCT Dong et al. (2022) transforms entity coordinates and represents each relation as a rotation in complex space. Rotate4D Le et al. (2023) employs two distinct rotational transformations to align the head embedding with the tail embedding. DCNE Dong et al. (2024) maps entities to dual complex number space, using rotations in 2D space via dual complex number multiplication to represent relations. TransERR Li et al. (2024) encodes knowledge graphs by rotating the head and tail entities with their corresponding unit quaternions.
@@
A common feature of these models is that the plausibility of the triplets is evaluated by calculating the distance between the head entity and the tail entity after transformation. However, these models fail to consider information about entities within the semantic space, leading to performance degradation.
@@ 3.2 Quaternion Algebra
The quaternion extends complex number system to four dimensions. In $n$-dimensional quaternion space $\mathbb{Q}^{n}$, a quaternion $\mathbf{p}\in\mathbb{Q}^{n}$ consists of one real component and three imaginary components. It can be formalized as: $\mathbf{p}=a+b\mathbf{i}+c\mathbf{j}+d\mathbf{k}$, where $a,b,c,d\in\mathbb{R}^{\frac{n}{4}}$ are real numbers and $\mathbf{i,j,k}$ are imaginary units. The imaginary part satisfies the Hamilton’s rules Hamilton (1844): $\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=\mathbf{ijk}=-1$.
@@ 5 Experiments
Datasets. To verify the effectiveness and robustness our model, we conducted extensive experiments on four standard knowledge graph completion datasets including WN18RR Dettmers et al. (2018), FB15k-237 Toutanova and Chen (2015), WN18 Bordes et al. (2013) and FB15k Bordes et al. (2013). The WN18 and FB15k datasets are known to suffer from a data leakage problem, leading to models being easily inferred and consequently performing well on metrics. WN18RR and FB15k-237 were derived as subsets of WN18 and FB15k respectively. These datasets are designed to address data leakage concerns and thereby present a more realistic prediction task. The detailed statistics of the four standard datasets are shown in Appendix B.
@@
Baselines. To verify the effectiveness of our model, we compared DaBR with several powerful baseline models, including both well-known and recently proposed ones with outstanding results. We divide these models according to the scoring function as follows:
1) Semantic Matching: TuckER Balazevic et al. (2019), QuatE Zhang et al. (2019), DualE Cao et al. (2021), QuatRE Nguyen et al. (2022).
2) Geometric Distance: ATTH Chami et al. (2020), Rotate3D Gao et al. (2020), Trans4E Nayyeri et al. (2021), RotateCT Dong et al. (2022), Rotate4D Le et al. (2023), CompoundE Ge et al. (2023), HAQE Liang et al. (2024d), DCNE Dong et al. (2024), FHRE Liang et al. (2024b) and TransERR Li et al. (2024).
@@ 5.1 Main Results
The KGC results on WN18 and FB15k datasets, as shown in Table 2. The Table 2 illustrates our model superiority over any previous model on the FB15k dataset. On the WN18 dataset, our model achieves the best results on all metrics, except for the H@1 metric which achieves second place. In conclusion, our model not only achieves optimal results compared to semantic matching models, but also achieves competitive results compared to geometric distance models.
@@ 6.1 Ablation Analysis
We show the results of the ablation experiments in Table 3. From the table, we can obtain the following conclusions: 1) The rotation of the tail entity and distance-adaptation are important parts of our model. 2) When our model removed the tail rotation, the model (i.e., Variant I) still achieved the best results compared to the models in the Table 1 and Table 2. We attribute this to the fact that our model can measure both the semantics of entities and the embedding distance of entities. 3) When our model removed distance-adaptation, the model (i.e., Variant II) performance decreased dramatically on all datasets. It is worth noting that our model still achieves optimal results on most datasets compared to the semantic matching model on most datasets.

### v2

@@ Abstract
Quaternion contains one real part and three imaginary parts, which provided a more expressive hypercomplex space for learning knowledge graph. Existing quaternion embedding models measure the plausibility of a triplet through either semantic matching or geometric distance scoring functions. However, it appears that semantic matching diminishes the separability of entities, while the distance scoring function weakens the semantics of entities. To address this issue, we propose a novel quaternion knowledge graph embedding model. Our model combines semantic matching with the geometric distance of entities to better measure the plausibility of triplets. Specifically, in the quaternion space, we perform a right rotation on head entity and a reverse rotation on tail entity to learn rich semantic features. We then utilize distance-adaptive translations to learn geometric distance between entities. Furthermore, we provide mathematical proofs to demonstrate our model can handle complex logical relationships. Extensive experimental results and analyses show our model significantly outperforms previous models on well-known knowledge graph completion benchmark datasets. Our code is available at https://github.com/llqy123/DaBR.
@@ 1 Introduction
Knowledge graphs (KGs) Liang et al. (2024a) are powerful tools for representing valid factual triplets by capturing entities and their relationships in a graphical format. Owing to the well-structured of graphs, KGs are often used for various Natural Language Processing tasks, such as question answering Mendes et al. (2024); Faldu et al. (2024), entity alignment Wang et al. (2024a, b), KG-based recommendation Liang et al. (2024c) and KG enhanced Large Language Model Wen et al. (2024).
However, KGs are usually incomplete and the incompleteness limits their application. As an effective tool for predicting missing facts, knowledge graph completion (KGC) has received considerable attention from researchers. Typically, researchers transform KGC tasks into knowledge graph embeddings (KGEs). KGE refers to learning representations of entities and relations in a low-dimensional space while preserving the graph’s inherent structure and semantic properties. In this representation space, a scoring function can be defined to measure the plausibility of each triplet, where valid triplets should receive higher scores than these invalid ones.
@@
Quaternion contains one real part and three imaginary parts, which providing a more expressive space for learning embeddings of entities and relations. Rotation in the quaternion space is often used to model the KGs. For example, QuatE Zhang et al. (2019) learns semantic information about entities by treating relations as rotations from head entities to tail entities. TransERR Li et al. (2024) encodes the KG by rotating the head and tail entities with their corresponding unit quaternions. These models use either semantic matching or distance scoring functions to measure the plausibility of the triplet, respectively. However, it appears that semantic matching diminishes the separability of entities, while the distance scoring function weakens the semantics of entities. For example, we visualized the results for the same query in Figure 1  [^1] For more information about queries, see Section 6.4.. Specifically, as shown in Figure 1, we observe that QuatE model overlaps some queries when using semantic matching as a scoring function. The entities of TransERR using the distance scoring function are also indistinguishable from each query.
To address this issue, we propose a Distance-adaptive quaternion knowledge graph embedding with Bidirectional Rotation model, named as DaBR. Our model combines semantic matching with the geometric distance of entities to better measure the plausibility of triplets. Specifically, in the quaternion space, we perform a right rotation on the head entity and a reverse rotation on the tail entity to learn rich semantic features. This process is called bidirectional rotation. We conducted extensive experiments on multiple well-known benchmark datasets for knowledge graph completion task. The experimental results and analyses demonstrated the effectiveness and robustness of our model.
Our contributions are summarized as follows:
  • We propose performing a right rotation on the head entity and a reverse rotation on the tail entity to learn rich semantic features.
  • We propose learning the embedding distance between entities by incorporating distance adaptive translations.
@@
  • Extensive experiments show that our model provides consistent and significant improvements over previous models in most metrics.
@@ 2 Related Work
For KGE models, the design of the scoring function directly affects these models’ performance and effectiveness. Based on the calculation methods of scoring functions in previous models, KGE scoring functions can mainly be categorized into semantic matching- and geometric distance-based.
@@
Semantic matching. Semantic matching scoring functions capture the interactions between entities and relations through inner products on embedding vectors. The hypothesis is that entities connected by relations are close to each other in the semantic space. For example, QuatE Zhang et al. (2019) obtains semantic information about entities through the Hamiltonian rotation of the head entity on the relation in quaternion space. DualE Cao et al. (2021) further enhances QuatE to model knowledge graphs in dual quaternion space. QuatRE Nguyen et al. (2022) associates each relation with two relation-aware rotations, which are used to rotate the quaternion embeddings of the head and tail entities, respectively. ConvQE Liang et al. (2024d) investigates the potential of quaternion convolution in knowledge graph embedding.
@@
A common feature of these models is the computation of the inner product between the head entity and the tail entity after a relation transformation. However, these models overlook the geometric distance properties between entities in the knowledge graph, which leads to distorted embeddings of the learned entities.
@@
Geometric distance. Geometric distance scoring functions assess the plausibility of triplets by calculating the distances between embedding vectors in the representation space. The goal of this scoring function is to keep the head/tail entity vector closer to the tail/head entity vector after being transformed through the relation vector. For example, TransE Bordes et al. (2013), considered the first model to employ a geometric distance scoring function, assumes that triplets $(h,r,t)$ in knowledge graphs should satisfy the expression $h+r\approx t$. However, TransE struggles with more complex relation types, such as one-to-many (1-to-N), many-to-one (N-to-1) and many-to-many (N-to-N).
@@
To address this limitation, several models using distance-based scoring functions have been proposed. For example, Rotate3D Gao et al. (2020) maps entities to a 3D space, defining the relation as a rotation from the head entity to the tail entity. Trans4E Nayyeri et al. (2021) performs rotations and translations in a quaternion space. RotateCT Dong et al. (2022) transforms entity coordinates and represents each relation as a rotation in complex space. Rotate4D Le et al. (2023) employs two distinct rotational transformations to align the head embedding with the tail embedding. DCNE Dong et al. (2024) maps entities to the dual complex number space, using rotations in the 2D space through the multiplication of dual complex numbers to represent relations. TransERR Li et al. (2024) encodes knowledge graphs by rotating the head and tail entities with their corresponding unit quaternions.
@@
A common feature of these models is that the plausibility of the triplets is evaluated by calculating the distance between the head entity and the tail entity after transformation. However, these models do not consider information about entities within the semantic space, leading to performance degradation.
@@ 3.2 Quaternion Algebra
The quaternion extends the complex number system to four dimensions. In $n$-dimensional quaternion space $\mathbb{Q}^{n}$, a quaternion $\mathbf{p}\in\mathbb{Q}^{n}$ consists of one real component and three imaginary components. It can be formalized as: $\mathbf{p}=a+b\mathbf{i}+c\mathbf{j}+d\mathbf{k}$, where $a,b,c,d\in\mathbb{R}^{\frac{n}{4}}$ are real numbers and $\mathbf{i,j,k}$ are imaginary units. The imaginary part satisfies the Hamilton’s rules Hamilton (1844): $\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=\mathbf{ijk}=-1$.
@@ 5 Experiments
Datasets. To verify the effectiveness and robustness of our model, we conducted extensive experiments on four standard knowledge graph completion datasets including WN18RR Dettmers et al. (2018), FB15k-237 Toutanova and Chen (2015), WN18 Bordes et al. (2013) and FB15k Bordes et al. (2013). The WN18 and FB15k datasets are known to suffer from a data leakage problem, which causes models to easily inferred and consequently performing well on metrics. WN18RR and FB15k-237 were derived as subsets of WN18 and FB15k respectively. These datasets are designed to address data leakage concerns and thereby present a more realistic prediction task. The detailed statistics of the four standard datasets are shown in Appendix B.
@@
Baselines. To verify the effectiveness of our model, we compared DaBR with several powerful baseline models, including both well-known and recently proposed ones with outstanding results. We divide these models according to the scoring function:
1) Semantic Matching: TuckER Balazevic et al. (2019), QuatE Zhang et al. (2019), DualE Cao et al. (2021), QuatRE Nguyen et al. (2022) and ConvQE Liang et al. (2024d).
2) Geometric Distance: ATTH Chami et al. (2020), Rotate3D Gao et al. (2020), Trans4E Nayyeri et al. (2021), RotateCT Dong et al. (2022), Rotate4D Le et al. (2023), CompoundE Ge et al. (2023), HAQE Liang et al. (2024e), DCNE Dong et al. (2024), FHRE Liang et al. (2024b) and TransERR Li et al. (2024).
@@ 5.1 Main Results
The KGC results on WN18 and FB15k datasets are shown in Table 2. The Table 2 illustrates our model superiority over any previous model on the FB15k dataset. On the WN18 dataset, our model achieves the best results on all metrics, except for the H@1 metric which achieves second place. In conclusion, our model not only achieves optimal results compared to semantic matching models, but also achieves competitive results compared to geometric distance models.
@@ 6.1 Ablation Analysis
We show the results of the ablation experiments in Table 3. From the table, we can obtain the following conclusions: 1) The rotation of the tail entity and distance-adaptation are important parts of our model. 2) When our model removed the tail rotation, the model (i.e., Variant I) still achieved the best results compared to the models in Table 1 and Table 2. We attribute this to the fact that our model can measure both the semantics of entities and the embedding distance of entities. 3) When our model removed distance-adaptation, the model (i.e., Variant II) performance decreased dramatically on all datasets. It is worth noting that our model still achieves optimal results on most datasets compared to the semantic matching model on most datasets.